OpenStudy (anonymous):

how much energy is needed to raise 32.6 g of water from 40c to 300c please explain

OpenStudy (aaronq):

apply this equation: \( q=m*C_p*\Delta T\) q= heat (energy) m=mass C_p = specific heat capacity of water (4.18 J/g*celsius) \(\Delta T=T_f-T_i\)= change in temperature